A cylindrical container of height equal to twice the diameter of its base can hold 12 liters (1L= 1,000 cm3) of water. Another cylindrical container with the same capacity has its height equal to three times the diameter of its base.

______________________ What is the diameter of the first container in cm?

_______________________ What is the diameter of the second container in cm?
1
Expert's answer
2020-12-04T12:03:34-0500

The volume of a cylinder is "V = \pi r^2 h" .

Knowing that "r=\frac{d}{2}" we can write the formula as "V = \pi (\frac{d}{2})^2 h" .

The following statements are given:

"V_1=V_2=12 * 1000 cm^3"

"h_1=2d_1"

"h_2=3d_2"

We can write the following equation for the first cylinder:

"V_1=\pi (\frac{d_1}{2})^2 h_1"

"12 * 1000cm^3=\pi (\frac{d_1}{2})^2 2d_1"

"12000cm^3=\pi \frac{(d_1)^3}{2}"

"d_1 = \sqrt[3]{\frac{12000cm^3*2}{\pi}}"

"d_1 \approx 19.695 cm"

We can write the following equation for the second cylinder:

"V_2=\pi (\frac{d_2}{2})^2 h_2"

"12 * 1000cm^3=\pi (\frac{d_2}{2})^2 3d_2"

"12000cm^3=\pi \frac{3(d_2)^3}{4}"

"d_2 = \sqrt[3]{\frac{12000cm^3*4}{3\pi}}"

"d_2 \approx 17.205 cm"

The answers to the given parts of the question are "19.695" and "17.205", respectively.